If a rectangle has a perimeter of 54 ft its sides are 9ft and has d ft
Answer by checkley79(3341)

P=2L+2WW
L=9
54=2*9+2W
54=18+2W
2W=54-18
2W=36
W=36/2
W=18 ANS. FOR THE WIDTH.
PROOF:
54=2*18+2*9
54=36+18
54=54